HUBBLE SPACE TELESCOPE DAILY REPORT #3078 PERIOD COVERED: 0000Z (UTC) 03/21/02 - 0000Z (UTC) 03/22/02 Daily Status Report as of 081/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed Fifteen Sets of NICMOS/1/2/3 8945 (Cooling System Monitoring) The Near Infrared Camera and Multi-Object Camera (NIC1, NIC2 and NIC3 was used 1}.to measure NICMOS detector performance during the cool-down and steady state operation of the NCS. 2}. to demonstrate stability {+/-0.1K} of the NICMOS detector temperature at the optimal science operating temperature. and 3}.to demonstrate repeatability {+/-0.1 K} of NICMOS detector temperature following changes from the optimal science operating temperature. No problems were encountered. 1.2 Completed Four Sets of ACS/WFC/HRC 8947 (Weekly Test) The Advanced Camera for Surveys (WFC and HRC) was used to perform basic tests to monitor, the read noise, the development of hot pixels and test for any source of noise in ACS CCD detectors. This program will be executed at least once a day for the entire lifetime of ACS. There were no problems reported. 1.3 Completed Two Sets of STIS/CCD 8964 (CCD Darks) The Space Telescope Imaging Spectrograph (CCD) was used to obtain dark frames twice daily, as in a corresponding calibration program. Standard calibration reference darks are constructed and hot pixel lists are updated for use in pipeline reduction of SMOV3B data. All observations completed without incident. 1.4 Completed Two Sets of STIS/CCD 8963 (CCD Bias) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD read noise from multiple bias frames obtained in common science modes: amplifier D, gains 1 and 4 electrons/DN, and binning factors of 1x1, 1x2, 2x1, and 2x2. The read noise measurements before and after NCC turn-on are then compared, and standard calibration reference files for use in pipeline reduction of SMOV3B data are constructed. The proposal had no reported problems. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 2 Successful: 2 Scheduled Re-acquisitions: 0 Successful: 0 2.2 FHST Updates: Scheduled: 5 Successful: 5 2.3 Operations Notes: The NCC cooldown continues nominally. Operations notes were used to: Turn on the ground limits for the FHST temperatures. Restore the FHST-2 mounting bracket temperature limits. Load a telemetry diagnostic. Per an operations request, the ACR3 limit check was enabled at 080/1322Z. The FHSTs were nominally powered on at 080/1457Z. The first map at 080/1626Z measured an error of -3140, 3072, -324 arcsecs (total RSS of 1.22 deg which agreed with previous Coarse Attitude Determination of 1.29 deg within the CAD error). The second map at 080/1802Z was used to produce an gyro bias update as well as the ARU/PRT to correct the attitude error. The gyro bias correction was 143, -155 and 68 arcsec/hr. The third map at 19:38 measured errors of -184, 204 and -81 arcsec. The fourth map at 080/2120Z was used to produce a second gyro bias update as well as a second ARU/PRT to correct the residual errors. The second gyro bias correction was 7, -4 and 6 arcsec/hr. The fifth map at 080/2252Z measured an error of -3, 0, 9 arcsec. The FHST update at 080/2321Z occurred in LOS after the type 2 slew at 080/2318Z. No FHST updates were scheduled prior to the slew. Correction values were not available at the time of this report issuance due to the LOS. The FHST update at 080/2325Z corrected an error of -1, 1 and -1 arcsec as seen by the telemetry at AOS at 080/2333Z. The FGS acquisition (1, 2, 2) (Search Radius of 90 arcsec) at 080/2331Z was successful and had initial #44 errors of 7.28 and 15.06 arcsecs in V2 and V3. Flight Operations was given the direction to perform an OBSINT to update the gyro triplet (3 gyro fail-down) bias matrices. The autonomous bias update process was enabled at the end of SM3B and will begin autonomously updating the biases for the 4 gyro configuration. ROP SR-3 was used to set the SSR-3 playback pointer at 080/1938Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of Servicing Mission Orbital Verification and the gradual resumption of normal science observations and calibrations. /CAW